The fourth term of an AP is 4. Then the sum of the first 7 terms is

Grade:12

1 Answers

bharat bajaj IIT Delhi
9 years ago
t4= 4
a + 3d = 4
S7= 3.5 x [2a + 6d]
S7= 7 (2)(a+3d) = 7(2)(4)
S7= 28
Thanks & Regards
Bharat Bajaj
askiitian faculty
IIT Delhi